# On the moduli space of simple sheaves on singular K3 surfaces

### v1

@@ 1. Introduction
    The moduli space $\operatorname{PSpl}(r;L,c_{2})$ is smooth of dimension $22rc_{2}-(r-1)L^{2}-2r^{2}+$ 2 (Proposition 3.1);
@@
(1) The subspace ${\mathcal{L}}$ is half dimensional in $\operatorname{PSpl}(r;L,c_{2})$ if and only if $-~\widetilde{{\mathcal{L}}}$ is half dimensional in $\operatorname{PSpl}(r+v;L,c_{2})$;
(2) The subspace $-~\widetilde{{\mathcal{L}}}$ is isotropic (respectively, Lagrangian) if and only if ${\mathcal{L}}$ is isotropic (respectively, Lagrangian).
@@
Acknowledgement. The first author is thankful for hospitality at Universitat de Barcelona and Henri Poincaré Institute, as well as to Sissa for a sabbatical year, during which part of this research took place. The first author would like the participants at the conference in Nantes 2023, in particular Kieran O’Grady, for useful conversations.
@@ 2.2. Moduli of simple perfect sheaves on surfaces
Let $X$ be a projective variety. A coherent sheaf ${\mathcal{F}}$ on $X$ is called perfect if it is perfect as an element of $D^{b}(\operatorname{Coh}(X))$; equivalently, if it admits a finite free resolution by locally free sheaves of finite rank.
@@ 2.4. Ext groups and trace maps
Let $X$ be a Gorenstein projective variety. For perfect sheaves $F$ and $G$ on $X$ the natural pairing
@@
$*\alpha(G,F):\operatorname{Ext}^{n-i}(G,F\otimes\omega_{X})\to\operatorname{Ext}^{i}(F,G)^{*}$
@@ 3.1. Moduli of perfect simple sheaf on singular K3 surfaces
Let $X$ be a K3 surface, $L\in Pic(X)$ and $r,c_{2}$ integers with $r\geq 1$. Then, the moduli space $\operatorname{PSpl}(r;L,c_{2})$ is smooth of dimension $22rc_{2}-(r-1)L^{2}-2r^{2}+$ 2.
@@ 3.2. Symplectic structure on moduli of simple perfect sheaves on smoothable K3 surfaces
Let $p:=(b,F,\alpha)$ be a point in ${\mathcal{M}}$, with $b\in B$, $F\in\operatorname{Coh}({\mathcal{X}}_{b})$ and $\alpha:\det F\cong{\mathcal{L}}|_{{\mathcal{X}}_{b}}$ The the relative tangent and obstruction spaces to the morphism $f$ at $p$ are $\operatorname{Ext}^{1}_{{\mathcal{X}}_{b}}(F,F)_{0}$ and $\operatorname{Ext}^{2}_{{\mathcal{X}}_{b}}(F,F)_{0}$; the result follows by the argument in Proposition 3.1. ∎
@@
Since each fiber $\operatorname{Ext}^{i}(F,F)$ of $\operatorname{{{\mathcal{E}}}xt}^{i}_{p}({\mathcal{F}},{\mathcal{F}})$ to a point $(b,F,\alpha)$ of ${\mathcal{M}}$ is invariant under the $\mu_{r}$ action on $F$, it follows that the sheaves $\operatorname{{{\mathcal{E}}}xt}^{i}_{p}({\mathcal{F}},{\mathcal{F}})$ are canonicaly isomorphic to $\varepsilon^{*}\mathbb{E}^{i}$, where $\mathbb{E}^{i}:=\varepsilon_{*}\operatorname{{{\mathcal{E}}}xt}^{i}_{p}({\mathcal{F}},{\mathcal{F}})$ are locally free on $\operatorname{PSpl}_{\pi}(r;{\mathcal{L}},c_{2})$. In particular, $\mathbb{E}^{1}$ is isomorphic to the relative tangent bundle of $\operatorname{PSpl}_{\pi}(r;{\mathcal{L}},c_{2})$ over $B$ since $\varepsilon$ is étale.
@@ 4.2. Generalized syzygy bundles
$\dim{\mathcal{L}}=\frac{1}{2}\dim Ext^{1}(F,F)=\frac{1}{1}\dim\operatorname{PSpl}(r;L,c_{2})$
@@
$\dim{\mathcal{L}}^{\prime}=\frac{1}{2}\dim Ext^{1}(S,S)=\frac{1}{2}\dim\operatorname{PSpl}(w-r;-L,L^{2}-c_{2})$
@@ 5.2. Smoothability of polarized K3 surfaces with isolated singularities
[…]
@@ 5.3. The case of quartics in $\mathbb{P}^{3}$
    $\operatorname{Ext}^{1}({\mathcal{P}},{\mathcal{O}}_{X})\to H^{0}(T^{1}_{X})$ is surjective if and only if $H^{1}((\mathbb{P}^{3},{\mathcal{J}}(4))=0$ where ${\mathcal{J}}$ is the Jacobian ideal sheaf of $X$.

### v2

@@ 1. Introduction
    The moduli space $\operatorname{PSpl}(r;L,c_{2})$ is a smooth algebraic space of dimension $22rc_{2}-(r-1)L^{2}-2r^{2}+$ 2 (Proposition 3.1);
@@
(1) The subspace ${\mathcal{L}}$ is half dimensional in $\operatorname{PSpl}(r;L,c_{2})$ if and only if $\widetilde{{\mathcal{L}}}$ is half dimensional in $\operatorname{PSpl}(r+v;L,c_{2})$;
(2) The subspace $\widetilde{{\mathcal{L}}}$ is isotropic (respectively, Lagrangian) if and only if ${\mathcal{L}}$ is isotropic (respectively, Lagrangian).
@@
Acknowledgement. The first author is thankful for hospitality at Universitat de Barcelona and Henri Poincaré Institute, as well as to Sissa for a sabbatical year, during which part of this research took place. The first author would like to thank the participants at the conference in Nantes 2023, in particular Kieran O’Grady, for useful conversations.
@@ 2.2. Moduli of simple perfect sheaves on surfaces
Let $X$ be a projective variety. A coherent sheaf ${\mathcal{F}}$ on $X$ is called perfect if it is perfect as an element of $D^{b}(\operatorname{Coh}(X))$; equivalently, if it admits a finite resolution by locally free sheaves of finite rank.
@@ 2.4. Ext groups and trace maps
Let $X$ be a Gorenstein projective variety of dimension $n$. For perfect sheaves $F$ and $G$ on $X$ the natural pairing
@@
$\alpha(G,F):\operatorname{Ext}^{n-i}(G,F\otimes\omega_{X})\to\operatorname{Ext}^{i}(F,G)^{*}$
@@ 3.1. Moduli of perfect simple sheaf on singular K3 surfaces
Let $X$ be a K3 surface, $L\in Pic(X)$ and $r,c_{2}$ integers with $r\geq 1$. Then, the moduli space $\operatorname{PSpl}(r;L,c_{2})$ is a smooth algebraic space of dimension $22rc_{2}-(r-1)L^{2}-2r^{2}+$ 2.
@@ 3.2. Symplectic structure on moduli of simple perfect sheaves on smoothable K3 surfaces
Let $p:=(b,F,\alpha)$ be a point in ${\mathcal{M}}$, with $b\in B$, $F\in\operatorname{Coh}({\mathcal{X}}_{b})$ and $\alpha:\det F\cong{\mathcal{L}}|_{{\mathcal{X}}_{b}}$ The relative tangent and obstruction spaces to the morphism $f$ at $p$ are $\operatorname{Ext}^{1}_{{\mathcal{X}}_{b}}(F,F)_{0}$ and $\operatorname{Ext}^{2}_{{\mathcal{X}}_{b}}(F,F)_{0}$; the result follows by the argument in Proposition 3.1. ∎
@@
Since each fiber $\operatorname{Ext}^{i}(F,F)$ of $\operatorname{{{\mathcal{E}}}xt}^{i}_{p}({\mathcal{F}},{\mathcal{F}})$ at a point $(b,F,\alpha)$ of ${\mathcal{M}}$ is invariant under the $\mu_{r}$ action on $F$, it follows that the sheaves $\operatorname{{{\mathcal{E}}}xt}^{i}_{p}({\mathcal{F}},{\mathcal{F}})$ are canonicaly isomorphic to $\varepsilon^{*}\mathbb{E}^{i}$, where $\mathbb{E}^{i}:=\varepsilon_{*}\operatorname{{{\mathcal{E}}}xt}^{i}_{p}({\mathcal{F}},{\mathcal{F}})$ are locally free on $\operatorname{PSpl}_{\pi}(r;{\mathcal{L}},c_{2})$. In particular, $\mathbb{E}^{1}$ is isomorphic to the relative tangent bundle of $\operatorname{PSpl}_{\pi}(r;{\mathcal{L}},c_{2})$ over $B$ since $\varepsilon$ is étale.
@@ 4.2. Generalized syzygy bundles
$\dim{\mathcal{L}}=\frac{1}{2}\dim Ext^{1}(F,F)_{0}=\frac{1}{2}\dim\operatorname{PSpl}(r;L,c_{2})$
@@
$\dim{\mathcal{L}}^{\prime}=\frac{1}{2}\dim Ext^{1}(S,S)_{0}=\frac{1}{2}\dim\operatorname{PSpl}(w-r;-L,L^{2}-c_{2})$
@@ 5.2. Smoothability of polarized K3 surfaces with isolated singularities
00{0}00{0}${H^{1}({\mathcal{P}}^{\vee})}$${\operatorname{Ext}^{1}({\mathcal{P}},{\mathcal{O}}_{X})}$${H^{0}(T_{X}^{1})}$${W}$${H^{1}(T_{X})}$${\operatorname{Ext}^{1}(\Omega_{X},{\mathcal{O}}_{X})}$${H^{0}(T_{X}^{1})}$00{0}${H^{2}({\mathcal{O}}_{X})}$${H^{2}({\mathcal{O}}_{X})}$${H^{2}({\mathcal{P}}^{\vee})}$${\operatorname{Ext}^{2}({\mathcal{P}},{\mathcal{O}}_{X})}$00{0}00{0}00{0}
@@ 5.3. The case of quartics in $\mathbb{P}^{3}$
    $\operatorname{Ext}^{1}({\mathcal{P}},{\mathcal{O}}_{X})\to H^{0}(T^{1}_{X})$ is surjective if and only if $H^{1}(\mathbb{P}^{3},{\mathcal{J}}(4))=0$ where ${\mathcal{J}}$ is the Jacobian ideal sheaf of $X$.